# Can Machines Learn Robustly, Privately, and Efficiently?

### 22 Dec 2023 v1

@@ Abstract.
The success of machine learning (ML) applications relies on vast datasets and distributed architectures, which, as they grow, present challenges for ML. In real-world scenarios, where data often contains sensitive information, issues like data poisoning and hardware failures are common. Ensuring privacy and robustness is vital for the broad adoption of ML in public life. This paper examines the costs associated with achieving these objectives in distributed architectures.
We overview the meanings of privacy and robustness in distributed ML, and clarify how they can be achieved efficiently in isolation. However, we contend that the integration of these objectives entails a notable compromise in computational efficiency. We delve into this intricate balance, exploring the challenges and solutions for privacy, robustness, and computational efficiency in ML applications.
Byzantine robustness, federated learning, decentralized learning, differential privacy, cryptography
ccs: Computing methodologies Machine learningccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Distributed systems securityccs: Computing methodologies Distributed algorithms
@@ 1. Introduction
Traditionally, the realms of robustness and privacy have been addressed in isolation in distributed ML, with separate strategies and techniques employed to enhance each of these two aspects. In practical systems, it is of paramount importance to address these fundamental issues in tandem. We explain in this paper that the confluence of privacy and robustness presents however a formidable challenge. While individual, efficient solutions have been proposed and empirically examined for both robust and private distributed ML (Kairouz et al., 2021; Allouah et al., 2023a), the direct amalgamation of these techniques to simultaneously ensure privacy and robustness is far from straightforward. As we shall explain, the merger of conventional privacy and robustness strategies in the context of distributed ML introduces a pivotal third dimension: efficiency. This dimension embodies the computational costs entailed by the proposed solutions. More specifically, we argue that safeguarding desirable performance or utility, often referring to accuracy in the context of machine learning, against strong and inquisitive adversaries necessitates a substantial escalation in computational resources. Nevertheless, we also provide insights into promising, computationally efficient alternatives that can deliver satisfactory accuracy for private and robust ML, when contending with weaker adversaries in terms of privacy and robustness.
@@
A Case in Point: Income Census. To illustrate the complex interplay between robustness and privacy, consider the aggregation of annual incomes in a census. The objective here is to compute the average annual income of a population, under the dual constraints of ensuring (a) individual income privacy and (b) robustness against misreported data. In pursuit of privacy, individuals may introduce independent Gaussian noise to their true income figures, thereby obscuring these figures from curious entities, as in differential privacy (Dwork et al., 2014). Simultaneously, for the sake of robustness, the aggregation of reported incomes might be executed robustly, such as by calculating the median income (Yin et al., 2018), to mitigate the impact of grossly corrupted income data. Nonetheless, despite its computational efficiency, the introduction of noise to individual incomes inevitably leads to estimation errors (Steinke and Ullman, 2016; Allouah et al., 2023c). On the other hand, the robust aggregation, as exemplified by the median, diverges substantially from the true average when income distributions exhibit significant skew. A natural compromise for attaining both privacy and robustness is the robust aggregation of noisy incomes. Regrettably, this approach compounds inaccuracies, because the median does not attenuate individual noises, as averaging does, even within a large population. An appealing alternative surfaces in the form of homomorphic encryption, enabling the encryption of individual incomes instead of injecting noise. This approach permits the computation of the average income within the encrypted data domain, bypassing the need for preliminary decryption. While homomorphic averaging can be achieved efficiently (Zhang et al., 2020), it is not robust against corrupted income data. Furthermore, implementing non-linear robust aggregation methods, like the median, in a homomorphic manner presents substantial challenges and computational costs. This inefficiency becomes particularly pronounced when applied to high-dimensional inputs, a scenario commonly encountered in modern deep learning (Choffrut et al., 2023). This simple example serves as a microcosm illustrating the intricate relationship between privacy, robustness, and efficiency within the context of distributed ML.
This Paper. We first give a concise overview of the definitions of robustness and privacy in distributed ML, along with the techniques employed to attain these goals. We emphasize that privacy and robustness can be efficiently achieved independently, even against the strongest adversaries. We then shed light on the intricate three-way trade-off between privacy, robustness, and efficiency. Lastly, we explain that this trade-off can be mitigated for weaker threat models.
@@ 2.2. Threats to Distributed ML
Leakage refers to the ability of a curious adversary to deduce private information about the data held by the workers. This inquisitive (a.k.a., passive) adversary could be the workers, the server, or an external entity. For example, the curious adversary might attempt to discern the presence of specific data points in the training dataset by strategically querying the training model, as in membership inference attacks (Shokri et al., 2017). Other attacks may attempt to approximate the private data of the workers, as in data reconstruction attacks (Zhu et al., 2019; Carlini et al., 2021), where the adversary analyzes the gradients computed during training. For clarity, we exclude privacy threats originating from external entities, specifically those only capable of accessing the finalized published model. Our focus remains solely on threats arising from participants within the system during training time, encompassing other workers and, if present, the server. In other words, we consider the local model of privacy (Kasiviswanathan et al., 2011; Allouah et al., 2023c), which constitutes the strongest privacy threat model. Moreover, within the centralized and federated settings, it is common to assume the server to be honest-but-curious: it adheres to the protocol by correctly aggregating the received messages. Yet, the server may exploit the received information to compromise data privacy.
2.3. Trustworthy ML
@@ 3.2. Robustness Techniques
Among these defense strategies, robust aggregation stands out by offering robust protection against the strongest (i.e., Byzantine) adversaries with minimal assumptions. Robust aggregation methods have not only demonstrated strong convergence guarantees (Blanchard et al., 2017; Yin et al., 2018; Farhadkhani et al., 2022; Allouah et al., 2023a, b) but have also been validated empirically (Farhadkhani et al., 2022). While the other defense techniques may present cost-effective alternatives, they either assume a less potent adversary (as in data sanitization) or rely on assumptions that may not hold in practice, such as homogeneous data distribution or the presence of a trusted proxy dataset on the server. Lastly, it is essential to mention that when workers’ data exhibits extreme heterogeneity, there is an inherent robustness cost attributable to Byzantine actors, irrespective of the chosen defense mechanism (Karimireddy et al., 2022; Allouah et al., 2023b). Interestingly, the challenge of data heterogeneity can be tackled by adding pre-aggregation robust operations, e.g., NNM (Allouah et al., 2023a) and Bucketing (Karimireddy et al., 2022), aimed at reducing the effect of heterogeneity among honest workers.
@@ 4. Trade-off between Privacy, Robustness, and Efficiency
The challenge of upholding privacy and robustness in distributed ML inherently introduces computational efficiency as a crucial third dimension. When confronted with the stringent demands of DP or cryptographic security, in conjunction with Byzantine robustness, the pursuit of both privacy and robustness does not seamlessly integrate with computational efficiency. We argue that the academic community must invest more research effort in this domain to yield efficient solutions that can maintain satisfactory accuracy. In alignment with this objective, we highlight promising research directions that involve relaxing privacy concerns and/or mitigating the Byzantine adversary, potentially paving the way for a more computationally feasible convergence of privacy and robustness in distributed ML.
@@ 4.1. Efficiently Achieving Privacy or Robustness Separately
Privacy and robustness can be guaranteed separately with satisfactory computational overhead. Indeed, privacy and efficiency are simultaneously achievable by the means of noise injection. Ensuring DP via noise injection is computationally cheap, as it only requires the addition of random noise for each model parameter, which is linear in the size of the model $d$. Moreover, it induces a non-negligible privacy-accuracy trade-off. Specifically, for strongly convex distributed learning problems involving $n$ workers with $m$ data points per worker, this trade-off can be quantified as a training loss error in the order of $\tfrac{d}{\varepsilon^{2}nm^{2}}$, where $\varepsilon$ is the example-level DP budget (Steinke and Ullman, 2016; Allouah et al., 2023c). However, this trade-off is effectively amortized in systems with a large number of workers $n$ or data points $m$.
@@ 4.2. The Computational Challenge of Privacy and Robustness
Differential Privacy and Robustness. A natural idea to efficiently ensure privacy and robustness is combining efficient techniques which guarantee these two objectives, when used in isolation. In fact, several works (Guerraoui et al., 2021; Zhu and Ling, 2022; Allouah et al., 2023c) have considered coupling noise injection for DP at the workers with robust aggregation methods at the server in the federated setting. However, Allouah et al. (2023c) have shown that there is a fundamental accuracy cost when simultaneously enforcing DP and Byzantine robustness. This accuracy cost predominantly arises from the inherent robustness cost associated with Byzantine workers in the presence of data heterogeneity (Karimireddy et al., 2022; Allouah et al., 2023b). In cases where the data across workers exhibits substantial heterogeneity, distinguishing the behavior of a minority of honest workers with outlier data from well-crafted malicious behavior becomes exceedingly challenging. Furthermore, broadly speaking, DP mechanisms introduce artificial data heterogeneity among workers through the injection of noise (Steinke and Ullman, 2016; Allouah et al., 2023c), thereby exacerbating the accuracy cost. Specifically, for strongly convex distributed learning problems involving $f$ Byzantine workers among $n$ workers, any Byzantine-robust and differentially private training algorithm must pay an additional training error scaling as $\tfrac{f}{n}\cdot\tfrac{1}{\varepsilon^{2}m^{2}}$ (Allouah et al., 2023c). In the worst case where there is a constant fraction of Byzantine workers in the system, this trade-off is not amortized even in systems with a large number of workers $n$, unlike the classical privacy-accuracy trade-off $\tfrac{d}{\varepsilon^{2}nm^{2}}$ (Allouah et al., 2023c). Yet, Allouah et al. (2023c) have introduced SMEA, a robust aggregation method that matches the aforementioned lower bound and attains the smallest training error among all known robust aggregation techniques. A noteworthy insight from SMEA is that it leverages the isotropic structure of the Gaussian noise used for DP to discard outliers. This contrasts with traditional robust aggregation methods, such as the coordinate-wise median and trimmed mean, which are not designed to work in conjunction with DP. However, SMEA is computationally inefficient, requiring an exponential runtime in the number of Byzantine workers (Allouah et al., 2023c). This emphasizes the complex trade-offs associated with combining DP and Byzantine-robust aggregation: one must choose between a traditional robust aggregation approach that offers reasonable execution times but comes with a considerable accuracy cost or SMEA, which theoretically provides tighter guarantees but incurs a considerable computational overhead.
@@
Cryptographic Security and Robustness. In contrast, HE provides an effective means to sidestep the lower bounds associated with DP (Steinke and Ullman, 2016; Allouah et al., 2023c). HE schemes offer the ability to efficiently perform fundamental mathematical operations, such as averaging (Zhang et al., 2020). However, when dealing with more complex non-linear operations, e.g., median, nearest neighbors, or min/max operators, which are pivotal components of well-known Byzantine-robust aggregation techniques (Blanchard et al., 2017; Yin et al., 2018; Allouah et al., 2023a), the homomorphic implementation incurs considerable computational overhead (Choffrut et al., 2023). Hence, the seamless integration of HE with robust aggregation methods demands meticulous calibration before deployment. This critical consideration has been tackled by Choffrut et al. (2023), who investigated the most suitable robust aggregation method for homomorphic implementation in terms of computational efficiency. Choffrut et al. (2023)’s analysis advocates for the suitability of the coordinate-wise trimmed mean and median aggregations (Yin et al., 2018). These methods exhibit a remarkable alignment with batch-friendly HE cryptosystems, harnessing the advantages of batch processing. This facilitates the simultaneous execution of multiple clear-text operations as a single homomorphic computation, resulting in a substantial reduction in computational cost. This compatibility is significantly important, particularly when dealing with large deep learning models with millions of parameters. Despite the computational gains introduced by batching, the homomorphic implementation of the trimmed mean and median still grapples with efficiency challenges. Choffrut et al. (2023) report that the time required for a single robust aggregation in the homomorphic domain exceeds 2 minutes, even when applied to a model featuring 713,000 parameters in a federated topology involving nine workers. This represents a 110-fold slowdown in comparison to the 1.1-second execution time of its non-robust counterpart (Zhang et al., 2020), despite the consideration of a larger model in the latter work. Furthermore, it is crucial to emphasize that ML algorithms typically run for several hundred iterations, amplifying the significant computational workload associated with this approach. These quantitative findings, which underscore the “impracticality” of even the most promising Byzantine-robust aggregations in terms of homomorphic cost, emphasize the great challenges faced in designing ML algorithms that must concurrently defend against the strongest privacy and robustness adversaries.
4.3. Weaker Threat Models
@@
So far, existing works have not been able to marry privacy, robustness, and efficiency against the strongest adversaries. Indeed, proposed solutions either suffer from a significant loss in accuracy via the use of DP (Zhu and Ling, 2022; Allouah et al., 2023c), or incur substantial computations when implementing robust aggregation methods using cryptographic primitives (Choffrut et al., 2023). Nevertheless, we argue that privacy and robustness can be efficiently combined when considering weaker adversaries. In general, in federated training, natural relaxations of the privacy threat model try to recover the honest non-curious server scenario, where the privacy-accuracy trade-off is known to be acceptable (Kairouz et al., 2021). Specifically, this trade-off is $n$ times smaller in theory, i.e., scaling as $\tfrac{d}{\varepsilon^{2}n^{2}m^{2}}$ for strongly convex learning problems, when trusting the server and protecting against a curious adversary only observing the global models (i.e., post aggregation at the server) during training. Furthermore, one can consider a federated-like setup with two non-colluding honest-but-curious servers to enable additive secret sharing techniques (Hao et al., 2021). These additional assumptions are also compatible with trust assumptions limiting the Byzantine attacks (as in suspicion-based defenses for instance), and enable more efficient solutions. For example, Hao et al. (2021) consider a federated-like setup where the two non-colluding servers possess proxy data for Byzantine detection. By doing so, they are able to train a model of $270,000$ parameters with $100$ workers in approximately $100$ seconds. The solution in (Hao et al., 2021) suffers a loss of at most $3\%$ in accuracy, when defending against a fraction of $20\%$ of Byzantine workers, compared to FedAvg (McMahan et al., 2017). One promising avenue for future exploration is the examination of the shuffle model (Cheu et al., 2019), which employs a trusted shuffler for worker identity anonymization, offering a relaxation of privacy threats with DP amplification guarantees. However, its study in the presence of Byzantine adversaries remains an unexplored research direction. Another interesting avenue for further research lies in weakening the conventional Byzantine threat model within the context of distributed ML. Instead of assuming that Byzantine workers consistently exhibit faults in every training iteration, one can consider an adversary who poisons workers’ models only with a certain probability (Yin et al., 2018). This weaker threat model aligns more closely with real-world practical systems and allows for the potential development of efficient HE-based solutions. These algorithms can capitalize on the existence of “optimistic” training iterations without faults, to sometimes perform homomorphic averaging instead of robust aggregation. This shift can significantly enhance the computational efficiency of the solution, at the cost of potentially weaker, probabilistic convergence guarantees.

### 11 Mar 2024 v2

# Robustness, Efficiency, or Privacy: Pick Two in Machine Learning

@@ Abstract.
The success of machine learning (ML) applications relies on vast datasets and distributed architectures which, as they grow, present major challenges. In real-world scenarios, where data often contains sensitive information, issues like data poisoning and hardware failures are common. Ensuring privacy and robustness is vital for the broad adoption of ML in public life. This paper examines the costs associated with achieving these objectives in distributed ML architectures, from both theoretical and empirical perspectives.
We overview the meanings of privacy and robustness in distributed ML, and clarify how they can be achieved efficiently in isolation. However, we contend that the integration of these two objectives entails a notable compromise in computational efficiency. In short, traditional noise injection hurts accuracy by concealing poisoned inputs, while cryptographic methods clash with poisoning defenses due to their non-linear nature. However, we outline future research directions aimed at reconciling this compromise with efficiency by considering weaker threat models.
Byzantine robustness, federated learning, decentralized machine learning, differential privacy
ccs: Computing methodologies Machine learningccs: Security and privacy Privacy-preserving protocolsccs: Computing methodologies Distributed algorithms
@@ 1. Introduction
Traditionally, the realms of robustness and privacy have been addressed in isolation in distributed ML, with separate strategies and techniques employed to enhance each of these two aspects. In practical systems, it is of paramount importance to address these fundamental issues in tandem. We explain in this paper that the confluence of privacy and robustness presents however a formidable challenge. While individual, efficient solutions have been proposed and empirically examined for both robust and private distributed ML (Kairouz et al., 2021; Allouah et al., 2023a), the direct amalgamation of these techniques to simultaneously ensure privacy and robustness is far from straightforward. As we shall explain, the merger of conventional privacy and robustness strategies in the context of distributed ML introduces a pivotal third dimension: efficiency. This dimension embodies the computational costs entailed by the proposed solutions. More specifically, we argue that safeguarding desirable performance or utility, often referring to accuracy in the context of machine learning, against strong and inquisitive adversaries necessitates a substantial escalation in computational resources. In essence, noise injection techniques, used to ensure privacy, result in a fundamental decrease in accuracy when combined with robustness primitives, whereas cryptographic methods do not combine efficiently with defenses against data poisoning due to their inherent requirement for non-linearity. Nevertheless, we also provide insights into promising, computationally efficient alternatives that can deliver satisfactory accuracy for private and robust ML, when contending with weaker adversaries in terms of privacy and robustness.
@@
A Case in Point: Income Census. To illustrate the complex interplay between robustness and privacy, consider the aggregation of annual incomes in a census. The objective here is to compute the average annual income of a population, under the dual constraints of ensuring (a) individual income privacy and (b) robustness against misreported data. In pursuit of privacy, individuals may introduce independent Gaussian noise to their true income figures, thereby obscuring these figures from curious entities, as in differential privacy (Dwork et al., 2014). Simultaneously, for the sake of robustness, the aggregation of reported incomes might be executed robustly, such as by calculating the median income (Yin et al., 2018), to mitigate the impact of grossly corrupted income data. Nonetheless, despite its computational efficiency, the introduction of noise to individual incomes inevitably leads to estimation errors (Steinke and Ullman, 2016; Allouah et al., 2023b). On the other hand, the robust aggregation, as exemplified by the median, diverges substantially from the true average when income distributions exhibit significant skew. A natural compromise for attaining both privacy and robustness is the robust aggregation of noisy incomes. Regrettably, this approach compounds inaccuracies, because the median does not attenuate individual noises, as averaging does, even within a large population. An appealing alternative surfaces in the form of homomorphic encryption, enabling the encryption of individual incomes instead of injecting noise. This approach permits the computation of the average income within the encrypted data domain, bypassing the need for preliminary decryption. While homomorphic averaging can be achieved efficiently (Zhang et al., 2020), it is not robust against corrupted income data. Furthermore, implementing non-linear robust aggregation methods, like the median, in a homomorphic manner presents substantial challenges and computational costs. This inefficiency becomes particularly pronounced when applied to high-dimensional inputs, a scenario commonly encountered in modern deep learning (Choffrut et al., 2023). This simple example serves as a microcosm illustrating the intricate relationship between privacy, robustness, and efficiency within the context of distributed ML.
Overview. The contributions of this paper are threefold. Firstly, we provide a concise overview of the definitions and techniques for robustness and privacy in distributed ML, aimed at a broad audience. Secondly, we introduce computational efficiency as a fundamental antagonist to the integration of privacy and robustness in distributed ML, exploring it through recent theoretical analyses and empirical observations for the first time. Lastly, we highlight future research directions for the academic community, where the three-way privacy-robustness-efficiency trade-off can be mitigated by focusing on weaker and more realistic threat models.
@@ 2.2. Threats to Distributed ML
Leakage refers to the ability of a curious adversary to deduce private information about the data held by the workers. This inquisitive (a.k.a., passive) adversary could be the workers, the server, or an external entity. For example, the curious adversary might attempt to discern the presence of specific data points in the training dataset by strategically querying the training model, as in membership inference attacks (Shokri et al., 2017). Other attacks may attempt to approximate the private data of the workers, as in data reconstruction attacks (Zhu et al., 2019; Carlini et al., 2021), where the adversary analyzes the gradients computed during training. For clarity, we exclude privacy threats originating from external entities, specifically those only capable of accessing the finalized published model. Our focus remains solely on threats arising from participants within the system during training time, encompassing other workers and, if present, the server. In other words, we consider the local model of privacy (Kasiviswanathan et al., 2011; Allouah et al., 2023b), which constitutes the strongest privacy threat model. Moreover, within the centralized and federated settings, it is common to assume the server to be honest-but-curious: it adheres to the protocol by correctly aggregating the received messages. Yet, the server may exploit the received information to compromise data privacy.
2.3. Formal Guarantees
@@ 3.2. Robustness Techniques
Among these defense strategies, robust aggregation stands out by offering robust protection against the strongest (i.e., Byzantine) adversaries with minimal assumptions. Robust aggregation methods have not only demonstrated strong convergence guarantees (Blanchard et al., 2017; Yin et al., 2018; Farhadkhani et al., 2022; Allouah et al., 2023a, 2024) but have also been validated empirically (Farhadkhani et al., 2022). While the other defense techniques may present cost-effective alternatives, they either assume a less potent adversary (as in data sanitization) or rely on assumptions that may not hold in practice, such as homogeneous data distribution or the presence of a trusted proxy dataset on the server. Lastly, it is essential to mention that when workers’ data exhibits extreme heterogeneity, there is an inherent robustness cost attributable to Byzantine actors, irrespective of the chosen defense mechanism (Karimireddy et al., 2022; Allouah et al., 2024). Interestingly, the challenge of data heterogeneity can be tackled by adding pre-aggregation robust operations, e.g., NNM (Allouah et al., 2023a) and Bucketing (Karimireddy et al., 2022), aimed at reducing the effect of heterogeneity among honest workers.
@@ 4. Trade-off between Privacy, Robustness, and Efficiency
The challenge of upholding privacy and robustness in distributed ML introduces computational efficiency as a crucial third dimension. When confronted with the requirements of DP or cryptographic security, in conjunction with Byzantine robustness, the pursuit of both privacy and robustness does not seamlessly integrate with computational efficiency.
@@ 4.1. Efficiently Achieving Privacy or Robustness Separately
Privacy and robustness can be guaranteed separately with satisfactory computational overhead. Indeed, privacy and efficiency are simultaneously achievable by the means of noise injection. Ensuring DP via noise injection is computationally cheap, as it only requires the addition of random noise for each model parameter, which is linear in the size of the model $d$. Moreover, it induces a non-negligible privacy-accuracy trade-off. Specifically, for strongly convex distributed learning problems involving $n$ workers with $m$ data points per worker, this trade-off can be quantified as a training loss error in the order of $\tfrac{d}{\varepsilon^{2}nm^{2}}$, where $\varepsilon$ is the example-level DP budget (Steinke and Ullman, 2016; Allouah et al., 2023b). However, this trade-off is effectively amortized in systems with a large number of workers $n$ or data points $m$.
@@ 4.2. The Computational Challenge of Privacy and Robustness
Differential Privacy and Robustness. A natural idea to efficiently ensure privacy and robustness is combining efficient techniques which guarantee these two objectives, when used in isolation. In fact, several works (Guerraoui et al., 2021; Zhu and Ling, 2022; Allouah et al., 2023b) have considered coupling noise injection for DP at the workers with robust aggregation methods at the server in the federated setting. However, Allouah et al. (2023b) have shown that there is a fundamental accuracy cost when simultaneously enforcing DP and Byzantine robustness. This accuracy cost predominantly arises from the inherent robustness cost associated with Byzantine workers in the presence of data heterogeneity (Karimireddy et al., 2022; Allouah et al., 2024). In cases where the data across workers exhibits substantial heterogeneity, distinguishing the behavior of a minority of honest workers with outlier data from well-crafted malicious behavior becomes exceedingly challenging. Furthermore, broadly speaking, DP mechanisms introduce artificial data heterogeneity among workers through the injection of noise (Steinke and Ullman, 2016; Allouah et al., 2023b), thereby exacerbating the accuracy cost. Specifically, for strongly convex distributed learning problems involving $f$ Byzantine workers among $n$ workers, any Byzantine-robust and differentially private training algorithm must pay an additional training error scaling as $\tfrac{f}{n}\cdot\tfrac{1}{\varepsilon^{2}m^{2}}$ (Allouah et al., 2023b). In the worst case where there is a constant fraction of Byzantine workers in the system, this trade-off is not amortized even in systems with a large number of workers $n$, unlike the classical privacy-accuracy trade-off $\tfrac{d}{\varepsilon^{2}nm^{2}}$ (Allouah et al., 2023b). Yet, Allouah et al. (2023b) have introduced SMEA, a robust aggregation method that matches the aforementioned lower bound and attains the smallest training error among all known robust aggregation techniques. A noteworthy insight from SMEA is that it leverages the isotropic structure of the Gaussian noise used for DP to discard outliers. This contrasts with traditional robust aggregation methods, such as the coordinate-wise median and trimmed mean, which are not designed to work in conjunction with DP. However, SMEA is computationally inefficient, requiring an exponential runtime in the number of Byzantine workers (Allouah et al., 2023b). This emphasizes the complex trade-offs associated with combining DP and Byzantine-robust aggregation: one must choose between a traditional robust aggregation approach that offers reasonable execution times but comes with a considerable accuracy cost or SMEA, which theoretically provides tighter guarantees but incurs a considerable computational overhead.
@@
Cryptographic Security and Robustness. In contrast, HE provides an effective means to sidestep the lower bounds associated with DP (Steinke and Ullman, 2016; Allouah et al., 2023b). HE schemes offer the ability to efficiently perform fundamental mathematical operations, such as averaging (Zhang et al., 2020). However, when dealing with more complex non-linear operations, e.g., median, nearest neighbors, or min/max operators, which are pivotal components of well-known Byzantine-robust aggregation techniques (Blanchard et al., 2017; Yin et al., 2018; Allouah et al., 2023a), the homomorphic implementation incurs considerable computational overhead (Choffrut et al., 2023). Hence, the seamless integration of HE with robust aggregation methods demands meticulous calibration before deployment. This critical consideration has been tackled by Choffrut et al. (2023), who investigated the most suitable robust aggregation method for homomorphic implementation in terms of computational efficiency. Choffrut et al. (2023)’s analysis advocates for the suitability of the coordinate-wise trimmed mean and median aggregations (Yin et al., 2018). These methods exhibit a remarkable alignment with batch-friendly HE cryptosystems, harnessing the advantages of batch processing. This facilitates the simultaneous execution of multiple clear-text operations as a single homomorphic computation, resulting in a substantial reduction in computational cost. This compatibility is significantly important, particularly when dealing with large deep learning models with millions of parameters. Despite the computational gains introduced by batching, the homomorphic implementation of the trimmed mean and median still grapples with efficiency challenges. Choffrut et al. (2023) report that the time required for a single robust aggregation in the homomorphic domain exceeds 2 minutes, even when applied to a model featuring 713,000 parameters in a federated topology involving nine workers. This represents a 110-fold slowdown in comparison to the 1.1-second execution time of its non-robust counterpart (Zhang et al., 2020), despite the consideration of a larger model in the latter work. Furthermore, it is crucial to emphasize that ML algorithms typically run for several hundred iterations, amplifying the significant computational workload associated with this approach. These quantitative findings, which underscore the “impracticality” of even the most promising Byzantine-robust aggregations in terms of homomorphic cost, emphasize the great challenges faced in designing ML algorithms that must concurrently defend against the strongest privacy and robustness adversaries.
5. Future Research Directions: Weaker Threat Models
@@
We conclude by highlighting promising research directions for the academic community, whereby adequately relaxing the privacy and robustness threat models could potentially improve utility. Indeed, proposed solutions either suffer from a significant loss in accuracy via the use of DP (Zhu and Ling, 2022; Allouah et al., 2023b), or incur substantial computations when implementing robust aggregation methods using cryptographic primitives (Choffrut et al., 2023). Nevertheless, we argue that privacy and robustness can be efficiently combined when considering weaker adversaries. In general, in federated training, natural relaxations of the privacy threat model try to recover the honest non-curious server scenario, where the privacy-accuracy trade-off is known to be acceptable (Kairouz et al., 2021). Specifically, this trade-off is $n$ times smaller in theory, i.e., scaling as $\tfrac{d}{\varepsilon^{2}n^{2}m^{2}}$ for strongly convex learning problems, when trusting the server and protecting against a curious adversary only observing the global models during training (i.e., post aggregation at the server). Furthermore, one can consider a federated-like setup with two non-colluding honest-but-curious servers to enable additive secret sharing techniques (Hao et al., 2021). These additional assumptions are also compatible with trust assumptions limiting the Byzantine attacks (as in suspicion-based defenses for instance), and enable more efficient solutions. For example, Hao et al. (2021) consider a federated-like setup where the two non-colluding servers possess proxy data for Byzantine detection. By doing so, they are able to train a model of $270,000$ parameters with $100$ workers in approximately $100$ seconds. The solution in (Hao et al., 2021) suffers a loss of at most $3\%$ in accuracy, when defending against a fraction of $20\%$ of Byzantine workers, compared to FedAvg (McMahan et al., 2017). One promising avenue for future work is the examination of the shuffle model (Cheu et al., 2019), which employs a trusted shuffler for worker identity anonymization, offering a relaxation of privacy threats with DP amplification guarantees. However, its study in the presence of Byzantine adversaries remains an unexplored research direction. Another interesting avenue for further research lies in weakening the conventional Byzantine threat model within the context of distributed ML. Instead of assuming that Byzantine workers consistently exhibit faults in every training iteration, one can consider an adversary who poisons workers’ models only with a certain probability (Yin et al., 2018). This weaker threat model aligns more closely with real-world practical systems and allows for the potential development of efficient HE-based solutions. These algorithms can capitalize on the existence of “optimistic” training iterations without faults, to sometimes perform homomorphic averaging instead of robust aggregation. This shift can significantly enhance the computational efficiency of the solution, at the cost of potentially weaker, probabilistic convergence guarantees.